# Node Perturbation Can Effectively Train Multi-Layer Neural Networks

### v6

@@ Abstract
Backpropagation (BP) remains the dominant and most successful method for training parameters of deep neural network models. However, BP relies on two computationally distinct phases, does not provide a satisfactory explanation of biological learning, and can be challenging to apply for training of networks with discontinuities or noisy node dynamics. By comparison, node perturbation (NP), also known as activity-perturbed forward gradients, proposes learning by the injection of noise into network activations, and subsequent measurement of the induced loss change. NP relies on two forward (inference) passes, does not make use of network derivatives, and has been proposed as a model for learning in biological systems. However, standard NP is highly data inefficient and can be unstable due to its unguided noise-based search process. In this work, we develop a number of alternative formulations of NP and relate them to the concept of directional derivatives as well as combining them with a decorrelating mechanism for layer-wise inputs. We find that a closer alignment with directional derivatives together with input decorrelation at every layer theoretically and practically enhances performance of NP learning with large improvements in parameter convergence and much higher performance on the test data, approaching that of BP. Furthermore, our novel formulation allows for application to noisy systems in which the noise process itself is inaccessible.
@@ 1 Introduction
Alternative algorithms have been put forth over the years, though their inability to scale and difficulty in achieving levels of performance comparable to BP have held back their use. Many focus either on greedy local learning rules or on random feedback matrices and projections, which are deemed more plausible than symmetric feedback weights (Dellaferrera and Kreiman, 2022; Frenkel et al., 2021; Kohan et al., 2018; Hinton, 2022)
The category of algorithms we focus on in this work are perturbation algorithms, such as node perturbation (NP) (Dembo and Kailath, 1990; Cauwenberghs, 1992), which require no feedback connectivity at all. In NP, node activations are perturbed by a small amount of random noise. Weights are then updated to produce the perturbed activations in proportion to the degree by which the perturbation improved performance. This method requires a measure of the loss function being optimized on a network both before and after the inclusion of noise. Such an approach is appealing because it leverages the same forward pass twice, rather than relying on two computationally distinct phases. It also does not require non-local information other than the global performance signal.
@@
In this work, we put forward three contributions. First, we reframe the process of node perturbation together with the subsequent measurement of the output loss change in terms of directional derivatives. Directional derivatives have been related to perturbation-based learning before in the context of forward gradient learning (Baydin et al., 2022; Ren et al., 2022) but never used to fully explicate a learning rule in this manner. This provides a more solid theoretical foundation for node perturbation on the whole.
Second, we introduce an approximation, referred to as activity-based node perturbation, which is more efficient and has the additional advantage that it can be implemented in noisy systems such as imprecise hardware implementations (Gokmen, 2021) or biological systems (Faisal et al., 2008), where the noise itself is not directly measurable.
Third, we propose to use a decorrelation method, first described by Ahmad et al. (2023), to de-bias activations. Because NP-style methods directly correlate perturbations of unit activities with changes in a reward signal, decorrelation of these unit activities helps to eliminate confounding effects, making credit assignment more straightforward. In addition, as demonstrated by Hiratani et al. (2022), correlations in the input data lead to more bias in NP’s updates by increasing their covariance. By combining decorrelation with the different NP methods, we find that it is possible to achieve orders of magnitude increase in model convergence speed, with performance levels rivaling networks trained by BP in certain contexts.
@@ 2.1 Node perturbation and its formulations
Let us define the forward pass of a fully-connected neural network, with $L$ layers, such that the output of a given layer, $l\in{1,2,\ldots,L}$ is given by $\mathbf{x}_{l}=f\left(\mathbf{a}_{l}\right)$, where $\mathbf{a}_{l}=\mathbf{W}_{l}\mathbf{x}_{l-1}$ is the pre-activation with weight matrix $\mathbf{W}_{l}$, $f$ is the activation function and $\mathbf{x}_{l}$ is the output from layer $l$. The input to our network is therefore denoted $\mathbf{x}_{0}$, and the output $\mathbf{x}_{L}$.
We consider learning rules which update the weights of such a network of the form
@@ 2.1.1 Iterative node perturbation
In the following, we develop a principled approach to node perturbation based learning.
Our goal is to determine the gradient of the loss with respect to the pre-activations in a layer $l$ by use of perturbations. To this end, we consider the partial derivative of the loss with respect to the pre-activation $\mathbf{a}^{i}_{l}$ of unit $i$ in layer $l$ for all $i$. We define a perturbed state as $\mathbf{\tilde{x}}_{k}(h)=f\left(\mathbf{\tilde{a}}_{k}+h\mathbf{m}_{k}\right)$
with $h$ an arbitrary scalar and binary vectors $\mathbf{m}_{k}=\mathbf{e}_{i}$ if $k=l$ with $\mathbf{e}_{i}$ a standard unit vector and $\mathbf{m}_{k}=\mathbf{0}$ otherwise. We may now define the partial derivatives as
@@
where $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ if $k=l$ and $\mathbf{v}_{k}=\mathbf{0}$ otherwise. Here, $\sigma^{2}$ is a scalar, equivalent to the noise variance.
@@
${\nabla}_{\mathbf{v}}\mathcal{L}=\lim_{h\rightarrow 0}\frac{\mathcal{L}\left(\mathbf{\tilde{x}}_{L}(h)\right)-\mathcal{L}\left(\mathbf{{x}}_{L}\right)}{h||\mathbf{v}||}\,$
where the directional derivative is measured by a difference in the loss induced in the vector direction $\mathbf{v}$, normalized by the vector length $||\mathbf{v}||$ and in the limit of infinitesimally small perturbation. The normalization ensures that the directional derivative is taken with respect to unit vectors. Note that this derivative is only being measured for layer $l$ as for all other layers the perturbation vector is composed of zeros.
As derived in Appendix B, by averaging the directional derivative across samples of the vector $\mathbf{v}$, in the limit $h\rightarrow 0$, we exactly recover the gradient of the loss with respect to a specific layer, such that
$\mathbf{g}_{l}=N_{l}\left\langle{\nabla}_{\mathbf{v}}\mathcal{L}\frac{\mathbf{v}_{l}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}\,.$ (2)
@@
$\Delta\mathbf{W}_{l}^{\textrm{INP}}=N_{l}\>\delta\mathcal{L}\frac{\mathbf{v}_{l}}{||\mathbf{v}||^{2}}\>\mathbf{x}_{l-1}^{\top}$ (3)
@@ 2.1.2 Node perturbation
Further zooming out, one may suppose that they wish to perturb all nodes in an entire network, rather even than just one layer at a time. In regular node perturbation, noise is injected into all layer’s pre-activations and weights are updated in the direction of the noise if the loss improves and in the opposite direction if it worsens. As with iterative node perturbation, two forward passes are required: one clean and one noise-perturbed. During the noisy pass, noise is injected into the pre-activation of each layer to yield a perturbed output
@@
As previously, one can measure a loss differential for an NP-based update. Supposing that the loss $\mathcal{L}$ is measured using the network outputs, the loss difference between the clean and noisy network is given by $\delta{\mathcal{L}}=\mathcal{L}(\mathbf{\tilde{x}}_{L})-\mathcal{L}(\mathbf{x}_{L})$, where $\delta{\mathcal{L}}$ is a scalar measure of the difference in loss induced by the addition of noise to the network. Given this loss difference and the network’s perturbed and unperturbed outputs, we compute a layer-wise learning signal by replacing the gradient $\mathbf{g}_{l}$ in Eq. 1 with a term consisting of the loss difference and a normalized noise vector:
@@
This update is similar to that of INP, with one key difference, that the normalization term, $\frac{N_{l}}{||v||^{2}_{2}}$, is approximated as the inverse of the node-wise variance, $\frac{1}{\sigma^{2}}$. Thus, regular NP is a network-wide measure of the directional derivative with an approximation of the required normalization.
@@ 2.1.3 Activity-based node perturbation
$\Delta\mathbf{W}_{l}^{\textrm{ANP}}=N\>\delta{\mathcal{L}}\frac{\delta{\mathbf{a}_{l}}}{||\delta\mathbf{{a}}||^{2}}\>\mathbf{x}_{l-1}^{\top}$ (6)
@@
The resulting update rule is referred to as activity-based node perturbation (ANP). Appendix C provides a derivation of this rule.
Here, we have now updated multiple aspects of the NP rule. First, rather than using the measure of isolated noise injected at each layer, we instead measure the change in activation between the clean and noisy networks. This has a theoretical benefit as well as a theoretical downside. As a theoretical benefit, by using the total noise at any given layer, this subsumes noise in all past layers and allows one to update a layer as if there was zero noise added to the preceeding layers. The only correlative noise issues remaining are then with the later network layers. ANP is thus demonstrably better than NP for layers closer to the output of a network.
On the other hand, the noise at each node in a layer is no longer independent and could now have a correlated structure. This is a downside in that it can skew the gradient descent direction, though can be shown to still converge to a minimum in loss. Both of these implications are expanded upon in Appendix C.
Second, this direct use of the activity difference also requires a recomputation of the scale of the perturbation vector and the requisite normalization. Unlike in the case of NP, one cannot assume the size of this normalization.
@@ 2.2 Increasing perturbation learning efficiency through decorrelation
Uncorrelated data variables have been proposed and demonstrated as impactful in making credit assignment more efficient in deep neural networks (LeCun et al., 2002). If a layer’s inputs, $\mathbf{x}_{l}$, have highly correlated features, a change in one feature can be associated with a change in another correlated feature, making it difficult for the network to disentangle the contributions of each feature to the loss. This can lead to less efficient learning, as has been described in previous research in the context of BP (Luo, 2017; Wadia et al., 2021). NP additionally benefits from decorrelation of input variables at every layer. Specifically,  Hiratani et al. (2022) demonstrate that the covariance of NP updates between layers $k$ and $l$ can be described as
$C^{\mathrm{np}}_{kl}\approx 2C^{\mathrm{sgd}}_{kl}+\delta_{kl}\left\langle\sum_{m=1}^{k}\|\mathbf{g}_{m}\|^{2}\mathbf{I}_{k}\otimes\mathbf{x}_{k-1}\mathbf{x}_{k-1}^{T}\right\rangle_{\mathbf{x}}$
where $C^{\mathrm{sgd}}_{kl}$ is the covariance of SGD updates, $\delta_{kl}$ is the Kronecker delta and $\otimes$ is a tensor product. The above equation implies that in NP the update covariance is twice that of the SGD updates plus an additional term that depends on the correlations in the input data $\mathbf{x}_{k-1}\mathbf{x}_{k-1}^{T}$. Removing correlations from the input data should therefore reduce the bias in the NP algorithm updates, possibly leading to better performance.

### v7

@@ Abstract
Backpropagation (BP) remains the dominant and most successful method for training parameters of deep neural network models. However, BP relies on two computationally distinct phases, does not provide a satisfactory explanation of biological learning, and can be challenging to apply for training of networks with discontinuities or noisy node dynamics. By comparison, node perturbation (NP), also known as activity-perturbed forward gradients, proposes learning by the injection of noise into network activations, and subsequent measurement of the induced loss change. NP relies on two forward (inference) passes, does not make use of network derivatives, and has been proposed as a model for learning in biological systems. However, standard NP is highly data inefficient and can be unstable due to its unguided noise-based search process. In this work, we develop a modern perspective on NP by relating it to the directional derivative and incorporating input decorrelation. We find that a closer alignment with directional derivatives together with input decorrelation at every layer theoretically and practically enhances performance of NP learning with large improvements in parameter convergence and much higher performance on the test data, approaching that of BP. Furthermore, our novel formulation allows for application to noisy systems in which the noise process itself is inaccessible, which is of particular interest for on-chip learning in neuromorphic systems.
@@ 1 Introduction
Alternative algorithms have been put forth over the years, though their inability to scale and difficulty in achieving levels of performance comparable to BP have held back their use. Many focus either on greedy local learning rules or on random feedback matrices and projections, which are deemed more plausible than symmetric feedback weights (Dellaferrera and Kreiman, 2022; Frenkel et al., 2021; Kohan et al., 2018; Hinton, 2022). The category of algorithms we focus on in this work are perturbation algorithms, specifically variants of node perturbation (NP) (Dembo and Kailath, 1990; Cauwenberghs, 1992), which require no feedback connectivity at all. In NP, node activations are perturbed by a small amount of random noise. Weights are then updated to produce the perturbed activations in proportion to the degree by which the perturbation improved performance. This method requires a measure of the loss function being optimized on a network both before and after the inclusion of noise. Such an approach is appealing because it leverages the same forward pass twice, rather than relying on two computationally distinct phases. It also does not require non-local information other than the global performance signal.
@@
In this work, we put forward three contributions. First, we reframe the process of node perturbation together with the subsequent measurement of the output loss change in terms of directional derivatives. Directional derivatives have been related to perturbation-based learning before in the context of forward gradient learning (Baydin et al., 2022; Ren et al., 2022) but never used to fully explicate a learning rule in this manner. This provides a more solid theoretical foundation for node perturbation on the whole. Second, we introduce an approximation, referred to as activity-based node perturbation, which is more efficient and has the additional advantage that it can be implemented in noisy systems such as imprecise hardware implementations (Gokmen, 2021) or biological systems (Faisal et al., 2008), where the noise itself is not directly measurable. Third, we propose to use a decorrelation method, first described by Ahmad et al. (2023), to de-bias activations. Because NP-style methods directly correlate perturbations of unit activities with changes in a reward signal, decorrelation of these unit activities helps to eliminate confounding effects. That is, it makes credit assignment more straightforward, effectively aligning gradient updates more closely with the natural gradient (Ahmad, 2024).
By combining decorrelation with the different NP methods, we find that it is possible to achieve orders of magnitude increase in model convergence speed, with performance levels rivaling networks trained by BP in certain contexts. Moreover, NP is a suitable candidate for learning on (noisy, resource-constrained) devices as it removes the need to implement the backward pass required for backpropagation. This is of particular interest for neuromorphic computing applications that make use of mixed-signal unconventional computing paradigms.
@@ 2.1 Node perturbation and its formulations
Let us define the forward pass of a fully-connected neural network, with $L$ layers, such that the output of a given layer, $l\in{1,2,\ldots,L}$ is given by $\mathbf{x}_{l}=f\left(\mathbf{a}_{l}\right)$, where $\mathbf{a}_{l}=\mathbf{W}_{l}\mathbf{x}_{l-1}$ is the pre-activation with weight matrix $\mathbf{W}_{l}$, $f$ is the activation function and $\mathbf{x}_{l}$ is the output from layer $l$. The input to our network is therefore denoted $\mathbf{x}_{0}$, and the output $\mathbf{x}_{L}$. We consider learning rules which update the weights of such a network of the form
@@ 2.1.1 Iterative node perturbation
In the following, we develop a principled approach to node perturbation based learning. Our goal is to determine the gradient of the loss with respect to the pre-activations in a layer $l$ by use of perturbations. To this end, we consider the partial derivative of the loss with respect to the pre-activation $\mathbf{a}^{i}_{l}$ of unit $i$ in layer $l$ for all $i$. We define a perturbed state as $\mathbf{\tilde{x}}_{k}(h)=f\left(\mathbf{\tilde{a}}_{k}+h\mathbf{m}_{k}\right)$ with $h$ an arbitrary scalar and binary vectors $\mathbf{m}_{k}=\mathbf{e}_{i}$ if $k=l$ with $\mathbf{e}_{i}$ a standard unit vector and $\mathbf{m}_{k}=\mathbf{0}$ otherwise.
We may now define the partial derivatives as
@@
where
$\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ if $k=l$ and $\mathbf{v}_{k}=\mathbf{0}$ otherwise. Here, $\sigma^{2}$ is a scalar, equivalent to the noise variance.
@@
[…]
where the directional derivative is measured by a difference in the loss induced in the vector direction $\mathbf{v}$, normalized by the vector length $\mathbf{v}||$ and in the limit of infinitesimally small perturbation. The normalization ensures that the directional derivative is taken with respect to unit vectors. Note that this derivative is only being measured for layer $l$ as for all other layers the perturbation vector is composed of zeros.
As derived in Appendix A, by averaging the directional derivative across samples of the vector $\mathbf{v}$, in the limit $h\rightarrow 0$, we exactly recover the gradient of the loss with respect to a specific layer, such that
[…]
@@
[…]
@@ 2.1.2 Node perturbation
Further zooming out, one may wish to perturb all nodes in an entire network, instead of just one layer at a time. In regular node perturbation, noise is injected into all layer’s pre-activations and weights are updated in the direction of the noise if the loss improves and in the opposite direction if it worsens. As with iterative node perturbation, two forward passes are required: one clean and one noise-perturbed. During the noisy pass, noise is injected into the pre-activation of each layer to yield a perturbed output
@@
As before, one can measure a loss differential for an NP-based update. Supposing that the loss $\mathcal{L}$ is measured using the network outputs, the loss difference between the clean and noisy network is given by $\delta{\mathcal{L}}=\mathcal{L}(\mathbf{\tilde{x}}_{L})-\mathcal{L}(\mathbf{x}_{L})$, where $\delta{\mathcal{L}}$ is a scalar measure of the difference in loss induced by the addition of noise to the network. Given this loss difference and the network’s perturbed and unperturbed outputs, we compute a layer-wise learning signal by replacing the gradient $\mathbf{g}_{l}$ in Eq. 1 with a term consisting of the loss difference and a normalized noise vector:
@@
[…]
@@ 2.1.3 Activity-based node perturbation
[…]
@@
The resulting update rule is referred to as activity-based node perturbation (ANP). Appendix B provides a derivation of this rule.
Here, we have now updated multiple aspects of the NP rule. First, rather than using the measure of isolated noise injected at each layer, we instead measure the change in activation between the clean and noisy networks. This has a theoretical benefit as well as a theoretical downside. As a theoretical benefit, by using the total noise at any given layer, this subsumes noise in all past layers and allows one to update a layer as if there was zero noise added to the preceding layers. The only correlative noise issues remaining are then with the later network layers. ANP is thus demonstrably better than NP for layers closer to the output of a network.
On the other hand, the noise at each node in a layer is no longer independent and could now have a correlated structure. This is a downside in that it can skew the gradient descent direction, though can be shown to still converge to a minimum in loss. Both of these implications are expanded upon in Appendix B. Second, this direct use of the activity difference also requires a recomputation of the scale of the perturbation vector and the requisite normalization. Unlike in the case of NP, one cannot assume the size of this normalization.
@@ 2.2 Increasing perturbation learning efficiency through decorrelation
Uncorrelated data variables have been proposed and demonstrated as impactful in making credit assignment more efficient in deep neural networks (LeCun et al., 2002). If a layer’s inputs, $\mathbf{x}_{l}$, have highly correlated features, a change in one feature can be associated with a change in another correlated feature, making it difficult for the network to disentangle the contributions of each feature to the loss. This can lead to less efficient learning, as has been described in previous research in the context of BP (Luo, 2017; Wadia et al., 2021). NP accordingly benefits from decorrelation of input variables at every layer. Specifically,  Hiratani et al. (2022) demonstrate that the covariance of NP updates between layers $k$ and $l$ can be described as
$C^{\mathrm{np}}_{kl}\approx 2C^{\mathrm{sgd}}_{kl}+\delta_{kl}\left\langle\sum_{m=1}^{k}\|\mathbf{g}_{m}\|^{2}\mathbf{I}_{k}\otimes\mathbf{x}_{k-1}\mathbf{x}_{k-1}^{T}\right\rangle_{\mathbf{x}}$ (7)
where $C^{\mathrm{sgd}}_{kl}$ is the covariance of SGD updates, $\delta_{kl}$ is the Kronecker delta and $\otimes$ is a tensor product. Equation 7 implies that in NP the update covariance is twice that of the SGD updates plus an additional term that depends on the correlations in the input data $\mathbf{x}_{k-1}\mathbf{x}_{k-1}^{T}$. Removing correlations from the input data should therefore reduce the bias in the NP algorithm updates, possibly leading to better performance.